### v1

@@ 4.3 Evaluation metrics
We used commonly used automatic text summarization evaluation metrics, including ROUGE [9], METEOR [10], MoverScore [11], BERTScore [12] and SciBERTScore [33], to assess the quality of generated title with the author-written title. ROUGE scores measure $n$-gram overlap between the generated and ground-truth titles. We use ROUGE-1, ROUGE-2, and ROUGE-L where the first uses unigram overlap, the second bigram overlap and the last compares the longest common subsequence between the generated title and the golden title. METEOR measures sentence-level accuracy based on the alignment between the generated text and the reference text. In contrast, MoverScore, BERTScore, and SciBERTScore aim to measure the semantic similarity between the output and the true title, using latent representations (embeddings) of the texts. MoverScore combines Word Mover’s Distance [34] and contextualized embeddings of the output and the ground-truth titles for semantic matching [11]. BERTScore calculates the cosine similarity between the BERT embeddings of the words in the two text sequences [12]. Since the current application is in the domain of computer science, we also develop a variant of BERTScore where we use SciBERT [33] to generate the embeddings, then we compute their cosine similarity as in BERTScore; we call the modified metric, SciBERTScore.
@@
These metrics are, however, inadequate to quantify factual consistency [35], which in our case connotes whether the entities present in the generated titles are precisely from the given abstract, and how much the entities in the generated title overlap with those in the author-written one. Hence, we have also used three new metrics, introduced by [13], to evaluate the factual consistency of the generated title.
@@
Let us call the author-written title as the target $t$, the model-generated title as the hypothesis $h$ and the input abstract as the source $s$. We define $\mathcal{N}(t)$ as the count of named-entities in the target (author-written title) and $\mathcal{N}(h)$ as the count of named-entities in the hypothesis (model-generated title). $\mathcal{N}(h\cap s)$ is the number of entity matches across the generated title and the input source document (abstract). Some named-entities in the title span multiple words. If even some words of a named entity in the title occur in the abstract, we consider it as a successful entity match. We used the scispaCy model en_core_sci_sm [^13] https://allenai.github.io/scispacy/ to identify entities. Precision-source, defined as $prec_{s}=\mathcal{N}(h\cap s)/\mathcal{N}(h)$, is a metric that is used to determine the intensity of hallucination in relation to the input source document (abstract). Note that $prec_{s}$ represents the percentage of entities mentioned in the generated title that can be retrieved from the input source document (abstract). Low $prec_{s}$ indicates the possibility of hallucination in the generated title. However, $prec_{s}$ does not capture the generated title’s entity-level correctness in relation to the author-written title. Therefore, recollecting that the author-written title is the target $t$, we define target-level entity accuracy in terms of precision-target $prec_{t}=\mathcal{N}(h\cap t)/\mathcal{N}(h);$ recall-target $recall_{t}=\mathcal{N}(h\cap t)/\mathcal{N}(t);$ and F1-target $F1_{t}=\frac{2*(recall_{t}*prec_{t})}{recall_{t}+prec_{t}}.$ Here, $\mathcal{N}(h\cap t)$ represents the number of matched named-entities in the generated title and the author-written title. Note that we have calculated the above mentioned precision-source and precision-target in two ways. One is to consider the entity mentioned in each document (which may be the source $s$ or target $t$ or hypothesis $h$) as a set so that multiple occurrences of an entity in a document are equivalent to a single occurrence. The other is to consider the entity mentioned in a input source (abstract) as a list; here, if a metric is defined as $\mu=\mathcal{N}(x\cap y)/\mathcal{N}(x)$, then for each entity mention in $x$, we check if it occurs in $y$, and if so, increment the intersection count $\mathcal{N}(x\cap y)$ by unity; this approach is followed in [13]. In the first approach, we denote the metrics as $prec_{s}^{U}$, $prec_{t}^{U}$, $recall_{t}^{U}$, and $F1_{t}^{U}$ ($U$ indicates that only unique entity mentions are considered). In the second, we represent them as $prec_{s}^{NU}$, $prec_{t}^{NU}$, $recall_{t}^{NU}$, and $F1_{t}^{NU}$ ($NU$ denoting not unique).
@@ 5.1 Quantitative comparison of different fine-tuned models
In this sub-section, we report the results of experiments on the test dataset of CSPubSum dataset as well as LREC-COLING-2024 dataset. Table 1 shows the performance in terms of ROUGE, METEOR, and semantic metrics like MoverScore, BERTScore and SciBERTScore, for the CSPubSum test dataset. Evaluation with entity-level factual consistency metrics precision-source($prec_{s}^{NU}$, $prec_{s}^{U}$), precision-target ($prec_{t}^{NU}$, $prec_{t}^{U}$), recall-target ($recall_{t}^{NU},recall_{t}^{U}$), and F1-target ($F1_{t}^{NU}$, $F1_{t}^{U}$) are shown in Table 2. The overall finding is that PEGASUS-large fine-tuned on the CSPubSum dataset achieves the highest scores for all the above metrics except precision-target metric. Table 3 and Table 4 show that the same PEGASUS-large model, that has been fine-tuned on the CSPubSum dataset, achieves the best performance on the LREC-COLING-2024 dataset except on BERTScore metric. Thus, the superlative performance of PEGASUS-large carries over to this dataset although it is not fine-tuned on it.
Table 1: Evaluation of all used models: F1-scores for ROUGE, METEOR are used for evaluating the both word-level matching and MoverScore, BERTScore, and SciBERTScore are used to compute the semantic fidelity of the generated titles for the CSPubSum dataset. All scores in percentage (%).
Model Name ROUGE-1 ROUGE-2 ROUGE-L METEOR MoverScore BERTScore SciBERTScore T5-base 44.25 25.04 38.92 38.36 38.09 89.9 76.06 BART-base 45.7 25.97 40.11 39.37 39.75 90.21 76.89 PEGASUS-large 46.75 27.13 40.67 42.61 40.43 90.35 76.93 LLaMA-3-8B 28.4 12.58 24.6 27.17 21.42 86.34 66.65 LLaMA-3-8B 40.8 21.23 36.57 34.5 37.02 89.99 76.41 ChatGPT-3.5 42.81 21.16 36.55 35.12 37.39 88.66 76.28
Table 2: Evaluation of all used models: precision-source, precision-target, recall-target and F1-target in terms of $prec_{s}^{NU}$, $prec_{s}^{U}$, $prec_{t}^{NU}$, $prec_{t}^{U}$, $recall_{t}^{NU}$, $recall_{t}^{U}$, $F1_{t}^{U}$ and $F1_{t}^{NU}$ scores are used for evaluating the factual consistency of the generated title for CSPubSum dataset. All scores in percentage (%).
Model Name $prec_{s}^{NU}$ $prec_{s}^{U}$ $prec_{t}^{NU}$ $recall_{t}^{NU}$ $F1_{t}^{NU}$ $prec_{t}^{U}$ $recall_{t}^{U}$ $F1_{t}^{U}$ T5-base 97.1 97.08 59.3 51.82 52.38 59.08 51.58 52.17 BART-base 97.44 97.39 61.52 52.84 53.91 61.35 52.56 53.72 PEGASUS-large 98.13 98.08 60.39 56.49 55.39 60.17 56.21 55.18 LLaMA-3-8B 67.7 67.7 39.19 40.48 36.81 38.96 39.98 36.48 LLaMA-3-8B 78.95 78.93 57.54 46.99 49.12 57.44 46.97 49.08 ChatGPT-3.5 92.73 92.73 59.79 50.95 51.7 59.79 51.01 51.74
Table 3: Evaluation of all used models: F1-scores for ROUGE, METEOR, MoverScore, BERTScore and SciBERTScore for LREC-COLING-2024 dataset. All scores in percentage (%).
Model Name ROUGE-1 ROUGE-2 ROUGE-L METEOR MoverScore BERTScore SciBERTScore T5-base 46.84 28.7 41.69 39.88 39.6 88.71 76.05 BART-base 46.87 27.66 41.89 38.93 39.61 88.84 75.94 PEGASUS-large 49.85 30.51 43.93 43.23 41.66 89.1 76.74 LLaMA-3-8B 32.92 16.66 27.66 30.61 25.68 86.77 67.42 LLaMA-3-8B 45.3 26.53 40.51 38.18 38.93 88.83 76.14 ChatGPT-3.5 45.16 23.97 38.88 37.45 38.85 89.54 75.64
Table 4: Evaluation of all used models: precision-source, precision-target, recall-target and F1-target in terms of $prec_{s}^{NU}$, $prec_{s}^{U}$, $prec_{t}^{NU}$, $prec_{t}^{U}$, $recall_{t}^{NU}$, $recall_{t}^{U}$, $F1_{t}^{U}$ and $F1_{t}^{NU}$ scores are used for evaluating the factual consistency of the generated title for LREC-COLING-2024 dataset. All scores in percentage (%).
Model Name $prec_{s}^{NU}$ $prec_{s}^{U}$ $prec_{t}^{NU}$ $recall_{t}^{NU}$ $F1_{t}^{NU}$ $prec_{t}^{U}$ $recall_{t}^{U}$ $F1_{t}^{U}$ T5-base 97.44 97.39 63.06 59.24 57.17 62.8 58.88 56.87 BART-base 96.24 96.23 64.3 59.21 57.78 64.11 58.92 57.56 PEGASUS-large 97.87 97.85 66.32 64.64 61.47 66.15 64.30 61.28 LLaMA-3-8B 74.6 74.52 45.29 48.16 43.48 45.02 47.64 43.13 LLaMA-3-8B 81.57 81.55 62.18 55.99 55.19 62.13 55.95 55.16 ChatGPT-3.5 91.98 91.98 60.87 57.66 55.56 60.84 57.69 55.56
@@
We believe the reason PEGASUS-large wins the competition is that it is trained with an objective function that favors summarization, i.e., generating important lines that have been masked in the input document. Thus, our observations show that although the LLMs show excellent performance on a wide range of NLP tasks, smaller models fine-tuned on custom datasets may exhibit better performance on a given task.
@@
A deep dive into the quantitative performance: Let us take a more nuanced view of the performance tables. Consider Table 1. It is immediately clear that the performance of LLaMA-3-8B (without fine-tuning) is considerably worse than that of the other models which are fine-tuned on the domain-specific data. This suggests that fine-tuning remains useful for attaining good model performance despite their extensive pre-training. In contrast, ChatGPT-3.5 performs better, even in a zero-shot setting, possibly because it is continuously used worldwide and as a result gets continually trained. A second observation is that although LLaMA-3-8B (fine-tuned) and ChatGPT-3.5 have smaller ROUGE and METEOR scores compared to the smaller pre-trained models, their performance is closer to that of the latter according to the semantic metrics like SciBERTScore; for example, SciBERTScore of all models except LLaMA-3-8B lie between 76.06 and 76.93. This indicates that all models (except LLaMA-3-8B) produce outputs with similar semantic similarity to the original titles. It also highlights the highly abstractive nature of the LLMs’ outputs, which, despite having low word overlap with the original titles, remain close in meaning to the original titles. Similar observations can be made from Table 3. Now let us analyze the entity overlap between the generated titles and the given titles and abstracts. Table 2 displays the figures for CSPubSum. First note that the precision-source as well as F1-target do not vary much whether or not we count multiple occurrences of an entity. Again, LLaMA-3-8B (without fine-tuning) shows the lowest scores for these metrics. The smaller pre-trained models perform similarly. Consider precision source $prec_{s}^{U}$: its large value indicates that the entities in the generated summary are mostly present in the input source which is the abstract here. $prec_{s}^{NU}=97.1$ for T5-base, $97.44$ for BART-base, and $98.13$ for PEGASUS-large. But $prec_{s}^{NU}$ is 92.73 for ChatGPT-3.5 and 78.95 for LLaMA-3-8B, showing that these models generate novel words more frequently. In contrast, the F1-target ($F1_{t}^{NU}$) lies in the 50’s for all models, meaning that there is only moderate overlap between the entities in the ground-truth title and the generated title. In particular, $F1_{t}^{NU}$ for LLMs is lower than that for the smaller models, again pointing to the abstractive nature of their output. We observe a similar pattern in Table 4 which reports the entity overlap statistics for the LREC-COLING-2024 dataset. Empirically, we did not observe noticeable hallucination in the generated titles but we intend to undertake a detailed study of this aspect in a future work.
@@ 5.2 Case studies
We show three representative examples of title generation one from CSPubSum (test subset) and two from LREC-COLING-2024 to illustrate the behavior of the models. Figure 1 shows the example from CSPubSum: the fine-tuned T5-base, BART-base, and PEGASUS-large models have generated titles that have significant similarity with the author-written titles. However, T5-base appears to generate a very long title. BART-base and PEGASUS-large generated the same titles. In contrast, LLaMA-3-8B without fine-tuning, produced a title that resembled a one-sentence summary and even that is incomplete; it is more of an extract from the abstract. After fine-tuning, the LLaMA-3-8B model generated an acceptable title. ChatGPT 3.5, using prompt-based techniques, successfully generated a title that also captures the essence of the paper and is similar to other generated titles and the author-written title. However, despite its stylistic flair, the evaluation metrics (in Table 1) indicate that the titles generated by ChatGPT-3.5 scored lower than those generated by the fine-tuned PEGASUS-large model. This discrepancy also highlights a limitation of the automated metrics in accurately evaluating titles generated in a highly abstractive and stylistic manner.
@@
fine-tuned T5-base: “Comparative statics effects independent of the utility function for portfolio choice and competitive firm under price uncertainty”
fine-tuned BART-base: “Comparative statics effects in the context of expected utility”
fine-tuned PEGASUS-large: “Comparative statics effects in the context of expected utility”
without fine-tuning LLaMA-3-8B: “The author proposes a methodological approach that enables comparative static analysis of various economic models, irrespective of the”
fine-tuned LLaMA-3-8B: “Comparative statics analysis: A new approach”
prompt based ChatGPT-3.5: “Comparative statics effects on expected utility in decision-making”
Figure 1: Input is an abstract from CSPubSum dataset. Title generated by each of the seven models is shown. Paper taken from https://www.sciencedirect.com/science/article/abs/pii/S037722171500586X
Now let us look at Fig. 2 which shows the outputs generated by the models for an example from LREC-COLING-2024 dataset. In case of BART-base, the generated title is incomplete. This is due to the hard limit on the output token count. We have observed this problem with other models, too. A simple workaround (without retraining the models) is to increase the output token limit during inference, which sometimes leads to a grammatically correct and complete title. While T5-base and fine-tuned LLaMA-3-8B generate a mostly correct title, they miss the “low-resource” setting which is correctly captured by PEGASUS-large. ChatGPT-3.5 generates almost the same title as that of PEGASUS-large except that it uses the abbreviation “NMT”, and abbreviations may not be desirable in a title.
@@
fine-tuned T5-base: “Reinforcement Learning Domain Adaptation for Neural Machine Translation”
fine-tuned BART-base: “A novel Reinforcement Learning Domain Adaptation method for Neural Machine Translation in the low-”
fine-tuned PEGASUS-large: “Reinforcement learning domain adaptation for Neural Machine Translation in the low-resource domain”
without fine-tuning LLaMA-3-8B: “Reinforced Domain Adaptation Method for Low Resource Neural Machine Translation”
fine-tuned LLaMA-3-8B: “Reinforcement learning domain adaptation for neural machine translation”
prompt based ChatGPT-3.5: “Reinforcement Learning Domain Adaptation for Low-Resource NMT”
Figure 2: Input is an abstract from LREC-COLING-2024 dataset. Title generated by each of the seven models is shown. Paper taken from https://aclanthology.org/2024.lrec-main.132/
In the example in Fig. 3, PEGASUS-large and fine-tuned LLaMA-3-8B have generated exactly the same title as the author-written one. A quick reading of the abstract of the input paper indicates that T5-base has captured an important aspect of the proposed model in the paper, namely the “entity abstraction approach”; however, the model has other characteristics and singling out one might not be appropriate. The title generated by BART-base is logically incomplete as it misses the part “for entailment tree generation”. LLaMA-3-8B without fine-tuning struggles to generate a succinct title. But ChatGPT-3.5 does generate an interesting title which captures the very purpose of the proposed model: “improving AI explanations”
@@
fine-tuned T5-base: “An entity abstraction approach for logical pattern memory pre-trained models”
fine-tuned BART-base: “Logical pattern memory pre-trained model”
fine-tuned PEGASUS-large: “A logical pattern memory pre-trained model for entailment tree generation”
without fine-tuning LLaMA-3-8B: “The proposed method addresses the limitations of previous approaches by incorporating an external memory structure to capture the latent representations”
fine-tuned LLaMA-3-8B: “Logical pattern memory pre-trained model for entailment tree generation”
prompt based ChatGPT-3.5: “Improving AI Explanations with Logical Pattern Memory Pre-trained Model”
Figure 3: Input is an abstract from LREC-COLING-2024 dataset. Title generated by each of the seven models is shown. Paper taken from https://aclanthology.org/2024.lrec-main.68/

### v2

@@ 4.3 Evaluation metrics
We used the commonly used automatic text summarization evaluation metrics, including ROUGE [9], METEOR [10], MoverScore [11], BERTScore [12] and SciBERTScore [38], to assess the quality of the generated titles with reference to the author-crafted titles. ROUGE scores measure $n$-gram overlap between the generated titles and the ground-truth titles. We have used ROUGE-1, ROUGE-2, and ROUGE-L where the first uses unigram overlap, the second bigram overlap, and the last compares the longest common subsequence between the generated title and the golden title. METEOR measures sentence-level accuracy based on the alignment between the generated text and the reference text. In contrast, MoverScore, BERTScore, and SciBERTScore aim to measure the semantic similarity between the output and the true title, using latent representations (embeddings) of the texts. MoverScore combines Word Mover’s Distance [39] and contextualized embeddings of the output and the ground-truth titles for semantic matching [11]. BERTScore calculates the cosine similarity between the BERT embeddings of the words in the two text sequences [12]. Since the current application is in the domain of computer science, we have developed a variant of BERTScore, where we used SciBERT [38] to generate the embeddings, then computed their cosine similarity as in BERTScore; we call the modified metric, SciBERTScore.
@@
However, these metrics are inadequate to quantify factual consistency [40], which, in our context, refers to whether the entities present in the generated titles also occur in the corresponding abstracts and the extent to which these entities overlap with those in the author-written titles. Therefore, we have utilized three new metrics, introduced by [13], to assess the factual consistency of the generated titles.
@@
Let us call the author-written title as the target $t$, the model-generated title as the hypothesis $h$ and the input abstract as the source $s$. We define $\mathcal{N}(t)$ as the count of named-entities in the target (author-written title) and $\mathcal{N}(h)$ as the count of named-entities in the hypothesis (model-generated title). Then, $\mathcal{N}(h\cap s)$ is the number of entity matches between the generated title and the input source (abstract). Some named-entities in the title span multiple words. If even some words of a named entity in the title occur in the abstract, we consider it as a successful entity match. We used the scispaCy model en_core_sci_sm [^16] https://allenai.github.io/scispacy/ to identify entities. Precision-source, defined as $prec_{s}=\mathcal{N}(h\cap s)/\mathcal{N}(h)$, is a metric used to assess the intensity of hallucination in relation to the input source (abstract). Note that $prec_{s}$ represents the percentage of entities mentioned in the generated title that can be retrieved from the input source document (abstract). Low $prec_{s}$ indicates the possibility of hallucination in the generated title. However, $prec_{s}$ does not capture the generated title’s entity-level correctness in relation to the author-written title. Therefore, recollecting that the author-written title is the target $t$, we define target-level entity accuracy in terms of precision-target $prec_{t}=\mathcal{N}(h\cap t)/\mathcal{N}(h);$ recall-target $recall_{t}=\mathcal{N}(h\cap t)/\mathcal{N}(t);$ and F1-target $F1_{t}=\frac{2*(recall_{t}*prec_{t})}{recall_{t}+prec_{t}}.$ Here, $\mathcal{N}(h\cap t)$ represents the number of matched named-entities in the generated title and the author-written title. Note that we have calculated the above mentioned precision-source and precision-target in two ways. The first considers entity mentions in each document (which may be the source $s$ or target $t$ or hypothesis $h$) as a set so that multiple occurrences of an entity in a document are equivalent to a single occurrence. The other treats the entity mentions in the input as a list. In this case, if a metric is defined as $\mu=\mathcal{N}(x\cap y)/\mathcal{N}(x)$, then for each entity mention in $x$, we check if it occurs in $y$; if it does, we increment the intersection count $\mathcal{N}(x\cap y)$ by unity; this approach is followed in [13]. In the first approach, we denote the metrics as $prec_{s}^{U}$, $prec_{t}^{U}$, $recall_{t}^{U}$, and $F1_{t}^{U}$ ($U$ indicates that only unique entity mentions are considered). In the second, we represent them as $prec_{s}^{NU}$, $prec_{t}^{NU}$, $recall_{t}^{NU}$, and $F1_{t}^{NU}$ ($NU$ denoting not unique).
@@ 5.1 Quantitative comparison of different fine-tuned models
In this sub-section, we report the results of experiments on the test dataset of CSPubSum dataset as well as LREC-COLING-2024 dataset. Table 1 shows the performance in terms of ROUGE, METEOR, and semantic metrics like MoverScore, BERTScore and SciBERTScore, for the CSPubSum test dataset. Evaluation with entity-level factual consistency metrics precision-source($prec_{s}^{NU}$, $prec_{s}^{U}$), precision-target ($prec_{t}^{NU}$, $prec_{t}^{U}$), recall-target ($recall_{t}^{NU},recall_{t}^{U}$), and F1-target ($F1_{t}^{NU}$, $F1_{t}^{U}$) are shown in Table 2.
The overall finding is that PEGASUS-large fine-tuned on the CSPubSum dataset achieves the highest scores for all the above metrics except the precision-target metric. Table 3 and Table 4 show that the same PEGASUS-large model, that has been fine-tuned on the CSPubSum dataset, achieves the best performance on the LREC-COLING-2024 dataset across all metrics except BERTScore. Thus, the superlative performance of PEGASUS-large in CSPubSum carries over to this dataset although it is not fine-tuned on it.
@@
We believe the reason PEGASUS-large wins the competition is that it is trained with an objective function that favors summarization, i.e., generating important lines that have been masked in the input document. Our observations show that although the LLMs show excellent performance on a wide range of NLP tasks, smaller models fine-tuned on custom datasets may exhibit better performance for a given task.
Table 1: Evaluation of the quality of the generated titles in CSPubSum.
Model Name ROUGE-1 ROUGE-2 ROUGE-L METEOR MoverScore BERTScore SciBERTScore T5-base 44.25 25.04 38.92 38.36 38.09 89.9 76.06 BART-base 45.7 25.97 40.11 39.37 39.75 90.21 76.89 PEGASUS-large 46.75 27.13 40.67 42.61 40.43 90.35 76.93 LLaMA-3-8B 28.4 12.58 24.6 27.17 21.42 86.34 66.65 LLaMA-3-8B 40.8 21.23 36.57 34.5 37.02 89.99 76.41 GPT-3.5-turbo 42.81 21.16 36.55 35.12 37.39 88.66 76.28
Table 2: Evaluation of factual consistency of the generated titles in CSPubSum.
Model Name $prec_{s}^{NU}$ $prec_{s}^{U}$ $prec_{t}^{NU}$ $recall_{t}^{NU}$ $F1_{t}^{NU}$ $prec_{t}^{U}$ $recall_{t}^{U}$ $F1_{t}^{U}$ T5-base 97.1 97.08 59.3 51.82 52.38 59.08 51.58 52.17 BART-base 97.44 97.39 61.52 52.84 53.91 61.35 52.56 53.72 PEGASUS-large 98.13 98.08 60.39 56.49 55.39 60.17 56.21 55.18 LLaMA-3-8B 67.7 67.7 39.19 40.48 36.81 38.96 39.98 36.48 LLaMA-3-8B 78.95 78.93 57.54 46.99 49.12 57.44 46.97 49.08 GPT-3.5-turbo 92.73 92.73 59.79 50.95 51.7 59.79 51.01 51.74
Table 3: Evaluation of the quality of the generated titles in LREC-COLING-2024.
Model Name ROUGE-1 ROUGE-2 ROUGE-L METEOR MoverScore BERTScore SciBERTScore T5-base 46.84 28.7 41.69 39.88 39.6 88.71 76.05 BART-base 46.87 27.66 41.89 38.93 39.61 88.84 75.94 PEGASUS-large 49.85 30.51 43.93 43.23 41.66 89.1 76.74 LLaMA-3-8B 32.92 16.66 27.66 30.61 25.68 86.77 67.42 LLaMA-3-8B 45.3 26.53 40.51 38.18 38.93 88.83 76.14 GPT-3.5-turbo 45.16 23.97 38.88 37.45 38.85 89.54 75.64
Table 4: Evaluation of factual consistency of the generated titles in LREC-COLING-2024.
Model Name $prec_{s}^{NU}$ $prec_{s}^{U}$ $prec_{t}^{NU}$ $recall_{t}^{NU}$ $F1_{t}^{NU}$ $prec_{t}^{U}$ $recall_{t}^{U}$ $F1_{t}^{U}$ T5-base 97.44 97.39 63.06 59.24 57.17 62.8 58.88 56.87 BART-base 96.24 96.23 64.3 59.21 57.78 64.11 58.92 57.56 PEGASUS-large 97.87 97.85 66.32 64.64 61.47 66.15 64.30 61.28 LLaMA-3-8B 74.6 74.52 45.29 48.16 43.48 45.02 47.64 43.13 LLaMA-3-8B 81.57 81.55 62.18 55.99 55.19 62.13 55.95 55.16 GPT-3.5-turbo 91.98 91.98 60.87 57.66 55.56 60.84 57.69 55.56
@@
Now, let us take a more nuanced view of the performance tables, beginning with the results on $n$-gram and semantic similarity between the author-written titles and the generated titles, as shown in Table 1. It is immediately clear that the performance of LLaMA-3-8B (without fine-tuning) is considerably worse than that of the other models which are fine-tuned on domain-specific data. This suggests that fine-tuning remains useful for attaining good performance despite the extensive pre-training that the models undergo. In contrast to LLaMA, GPT-3.5-turbo performs better, even in a zero-shot setting, possibly because it is continuously used worldwide and as a result gets continually trained. A second observation is that although LLaMA-3-8B (fine-tuned) and GPT-3.5-turbo have smaller ROUGE and METEOR scores compared to the smaller pre-trained models, their performance is closer to that of the latter according to the semantic metrics like SciBERTScore. For example, SciBERTScore values of all models except LLaMA-3-8B in Table 1 lie between 76.06 and 76.93. This indicates that all models (except LLaMA-3-8B) produce outputs with similar semantic similarity to the original titles. It also highlights the highly abstractive nature of the LLMs’ outputs, which exhibit low word overlap with the original titles yet remain semantically close to them. Similar observations can be made from Table 3.
Now let us analyze the entity overlap between the generated titles and the author-written titles and abstracts. Table 2 displays the figures for CSPubSum. First note that the precision-source as well as F1-target do not vary much whether or not we count multiple occurrences of an entity. Again, LLaMA-3-8B (without fine-tuning) achieves the lowest scores for these metrics. The smaller pre-trained models display similar performance to one another. Consider precision source $prec_{s}^{NU}$: its large value indicates that the entities in the generated summary are mostly present in the input source which is the abstract here. $prec_{s}^{NU}=97.1$ for T5-base, $97.44$ for BART-base, and $98.13$ for PEGASUS-large. But $prec_{s}^{NU}$ is 92.73 for GPT-3.5-turbo and 78.95 for LLaMA-3-8B, showing that these models generate novel words more frequently. In contrast, the F1-target ($F1_{t}^{NU}$) lies in the 50’s for all models, meaning that there is only moderate overlap between the entities in the ground-truth title and the generated title – thus, a generated titles does not match the golden title quite well but, at least for smaller models, it generally does not include entities outside the given abstract. Note that $F1_{t}^{NU}$ for LLMs is lower than that for the smaller models, again pointing to the abstractive nature of their output. We observe a similar pattern in Table 4 which reports the entity overlap statistics for the LREC-COLING-2024 dataset. Clearly, the lower values of $prec_{s}^{NU|U}$ and $F1_{t}^{NU|U}$ for LLMs could indicate hallucination. But upon manually examining a few generated titles we did not notice any instances of hallucination. This hints at the presence of novel words that preserve the intended meaning.
@@ 5.2 Case studies
We show three representative examples of title generation, one from CSPubSum (test subset) and two from LREC-COLING-2024 to illustrate the behavior of the models. Fig. 1 shows the example from CSPubSum: the fine-tuned T5-base, BART-base, and PEGASUS-large models have generated titles that have significant similarity with the author-written titles. However, T5-base has generated a very long title. BART-base and PEGASUS-large generated the same titles. In contrast, LLaMA-3-8B without fine-tuning, produced a title that resembled a one-sentence summary and even that is incomplete; it is more of an extract from the abstract. After fine-tuning, the LLaMA-3-8B model generated an acceptable title. GPT-3.5-turbo, using prompt-based techniques, successfully generated a title that also captures the essence of the paper and is similar to other generated titles and the author-written title. Recollect that according to the evaluation metrics in Table 1, the titles generated by GPT-3.5-turbo score lower than those generated by the fine-tuned PEGASUS-large model, despite the former’s stylistic flair. This discrepancy highlights a limitation of the automated metrics in accurately evaluating titles generated in a highly abstractive and stylistic manner.
@@
T5-base: “Comparative statics effects independent of the utility function for portfolio choice and competitive firm under price uncertainty”
BART-base: “Comparative statics effects in the context of expected utility”
PEGASUS-large: “Comparative statics effects in the context of expected utility”
LLaMA-3-8B: “The author proposes a methodological approach that enables comparative static analysis of various economic models, irrespective of the”
LLaMA-3-8B: “Comparative statics analysis: A new approach”
Zero-shot GPT-3.5-turbo: “Comparative statics effects on expected utility in decision-making”
Figure 1: Input is an abstract from CSPubSum. Titles generated by the different models are shown. Paper taken from https://www.sciencedirect.com/science/article/abs/pii/S037722171500586X.
Now let us look at Fig. 2, which shows the outputs generated by the models given an abstract from the LREC-COLING-2024 dataset. In case of BART-base, the generated title is incomplete. This is due to the hard limit on the output token count. We have observed this problem with other models, too. A simple workaround (without retraining the models) is to increase the output token limit during inference, which sometimes leads to a grammatically correct and semantically complete title. While T5-base and fine-tuned LLaMA-3-8B have generated acceptable titles, they missed the “low-resource” setting which is accurately captured by PEGASUS-large. GPT-3.5-turbo generated almost the same title as that of PEGASUS-large, except that it uses the abbreviation “NMT”, and abbreviations may not be desirable in a title.
@@
T5-base: “Reinforcement Learning Domain Adaptation for Neural Machine Translation”
BART-base: “A novel Reinforcement Learning Domain Adaptation method for Neural Machine Translation in the low-”
PEGASUS-large: “Reinforcement learning domain adaptation for Neural Machine Translation in the low-resource domain”
LLaMA-3-8B: “Reinforced Domain Adaptation Method for Low Resource Neural Machine Translation”
LLaMA-3-8B: “Reinforcement learning domain adaptation for neural machine translation”
Zero-shot GPT-3.5-turbo: “Reinforcement Learning Domain Adaptation for Low-Resource NMT”
Figure 2: Input is an abstract from LREC-COLING-2024. Titles generated by the different models are shown. Paper taken from https://aclanthology.org/2024.lrec-main.132/.
In the example shown in Fig. 3, PEGASUS-large and fine-tuned LLaMA-3-8B generated titles that are identical to the author-written title. A quick reading of the abstract of the input paper indicates that T5-base has captured an important aspect of the proposed model in the paper, namely the “entity abstraction approach”; however, the model in the paper has other characteristics and singling out one may not be appropriate. The title generated by BART-base is logically incomplete as it misses the phrase “for entailment tree generation”. LLaMA-3-8B without fine-tuning struggled to generate a succinct title. In contrast, GPT-3.5-turbo generated an interesting title which captures the very purpose of the proposed model: “improving AI explanations”.
@@
T5-base: “An entity abstraction approach for logical pattern memory pre-trained models”
BART-base: “Logical pattern memory pre-trained model”
PEGASUS-large: “A logical pattern memory pre-trained model for entailment tree generation”
LLaMA-3-8B: “The proposed method addresses the limitations of previous approaches by incorporating an external memory structure to capture the latent representations”
LLaMA-3-8B: “Logical pattern memory pre-trained model for entailment tree generation”
Zero-shot GPT-3.5-turbo: “Improving AI Explanations with Logical Pattern Memory Pre-trained Model”
Figure 3: Input is an abstract from LREC-COLING-2024. Titles generated by the different models are shown. Paper taken from https://aclanthology.org/2024.lrec-main.68/.